Factorise:x^4+x^2+1
1
Expert's answer
2012-06-12T10:11:15-0400
x^4+x^2+1=x^4+2*x^2+1-x^2=(x^4+2*x^2+1)-x^2=(x^2 + 1)^2 - ( x
)^2=[x^2+1+x]*[x^2+1-x]